OpenStudy (anonymous):

Factor the expression given below. Write each factor as a Polynomial in descending order. 4x^2-1

OpenStudy (anonymous):

To factor 4x^2 - 1 use difference of squares formula, because both terms are perfect squares. The formula is a^2 - b^2 = (a - b)(a + b) In this example we have 4x^2 - 1 = (2 x)^2 - 12 = (2 x - 1)(2 x + 1)

OpenStudy (anonymous):

Do you understanf @ObamaBinLlama ?

OpenStudy (anonymous):

Gimme a min. :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

(2x+1)(2x-1) ? ;D

OpenStudy (anonymous):

yes congrats :)

OpenStudy (anonymous):

May I have a medal? :)

OpenStudy (anonymous):

/Given Medal /Closed! ;D